a rectangle has sides with lengths of 6 and 9 find the lengths of the diagonals of the rectangle
Answer by checkley77(12844)

a^2+b^2=c^2
6^2+9^2=c^2
36+81=c^2
127=c^2
c=sqrt127
c=11.269 ans.